OpenStudy (anonymous):

Solve for x: |x − 2| + 10 = 12

OpenStudy (anonymous):

x = 4

OpenStudy (anonymous):

x = 0, x = 4

OpenStudy (anonymous):

bot of your answers are incorrect martin and spidey

OpenStudy (anonymous):

x-2+10 = 12 and x+2+10 = 12 X = 0 , x = 4

OpenStudy (anonymous):

and zjrod

OpenStudy (anonymous):

mine is right :)

OpenStudy (solomonzelman):

\(\normalsize\color{blue}{ |x − 2| + 10 = 12 }\) subtract 10 from both sides \(\normalsize\color{blue}{ |x − 2|+ 10 \color{red}{-10}= 12 \color{red}{-10}}\) \(\normalsize\color{blue}{ |x − 2| = 2 }\) \(\normalsize\color{blue}{ |x − 2| = 2 ~~~~\color{red}{⇒} ~~~~~~~~~~~~~~~~~~~~~~x-2=2,~~~~or~~~~~~x-2=-2}\)

OpenStudy (anonymous):

you guys forgot -4 and +4

OpenStudy (anonymous):

Now how about a medal for sidey

OpenStudy (solomonzelman):

@FriedRice , no you are incorrect.

OpenStudy (anonymous):

yes i am correct try plugging it in the equation

OpenStudy (anonymous):

you guys are slacking!!!

OpenStudy (anonymous):

-4 plugs in and equals 16 not 12

OpenStudy (anonymous):

oooops im slacking lol

OpenStudy (anonymous):

nvm that im soooooo sorry

OpenStudy (anonymous):

its ok :)

OpenStudy (anonymous):

lol